If f(x)=2x^3+Ax^2+5x-7 and f(2)=7, what is the value of A?
Answer by Timnewman(323)

hi dear,
just put 2 in f(x) and equate it to 7
i.e
2(2)^3+A(2)²+5(2)-7=7
simplify and get;
16+4A+10-7=7
4A+19=7
4A=7-19
4A=-12
A=-3
Hope this helps?